OpenStudy (anonymous):

A total of 308 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (anonymous):

@ksanka

OpenStudy (anonymous):

a - the number of adult's tickets then 3a - the number of student's tickets 3a+a=308 4a=308 a=308/4=77

OpenStudy (anonymous):

A garden table and a bench cost $592 combined. The garden table costs $58 less than the bench. What is the cost of the bench?

OpenStudy (anonymous):

@ksanka Thanks so much, I have a couple more and i'll be done

OpenStudy (anonymous):

ok. a - the cost of table a+58 - cost of bench a+a+58=592 2a+58=592

OpenStudy (anonymous):

a=267

OpenStudy (anonymous):

this is table, so bench costs: 267+58=325

OpenStudy (anonymous):

Alvin is 11 years older than Elga. The sum of their ages is 93 . What is Elga's age?

OpenStudy (anonymous):

@ksanka

OpenStudy (anonymous):

@ksanka